Database last updated on Thu, 05 Feb 2026 21:40:50 +0000 by localhost

Critical warnings from b-files' comments for days between 2022094 and 2022094 (0 sec, 7 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 093 05 36 32 Reset Warm -> Tepid from 2022 093 05 36 26 until 2022 093 05 36 27 . Counts: 1,0313B09422.194Toron#19405:26:251-1
rl:2022 093 05 38 57 Bad command string. Command string buffer cleared.303B09422.194Toron#19405:26:281-1
rl:2022 093 05 38 57 Bad command string. Command string buffer cleared.303B09422.194Toron#19405:26:281-1
rl:2022 093 05 38 57 <303B09422.194Toron#19405:26:301-1
rl:2022 094 17 25 50 Command not accepted while another in progress.11B09422.194Toron#19405:29:081-1
hp:Spectrometer alignment trouble.21B09422.194Toron#19416:28:121-1
ab_sk:Brewer failed to respond 5 times11B09422.194Toron#19418:03:261-1

Instrument constant changes for the week before 2022094 (5 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
1942022094616.5

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
1942022094230.2

Non-critical warnings from b-files' comments for days between 2022094 and 2022094 (0 sec, 11 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 093 05 36 32 Reset Warm -> Tepid from 2022 093 05 36 26 until 2022 093 05 36 27 . Counts: 1,013B09422.194Toron#19405:26:251-1
rl:2022 093 05 38 57 Bad command string. Command string buffer cleared.13B09422.194Toron#19405:26:281-1
rl:2022 093 05 38 57 Bad command string. Command string buffer cleared.13B09422.194Toron#19405:26:281-1
rl:2022 093 05 38 57 <13B09422.194Toron#19405:26:301-1
rl:2022 093 05 52 07 New value set for watermark SUPPLY.VOLTAGE.HIGH[-5VSEC]: -4.693947 Volts.13B09422.194Toron#19405:26:321-1
rl:2022 094 14 32 13 New value set for watermark SUPPLY.VOLTAGE.HIGH[-5VSEC]: -4.683193 Volts.33B09422.194Toron#19405:26:351-1
hg:HOME key pressed399B09422.194Toron#19411:06:481-1
hg:Going to menu399B09422.194Toron#19411:06:511-1
ds:Suppressed oscillation of filter wheel.15B09422.194Toron#19414:00:011-1
ds:DS intensity too high ( 219907/cy) quitting.199B09422.194Toron#19415:15:301-1
fr:MOTOR ORIGINS CORRUPTED. QUITTING.199B09422.194Toron#19418:43:371-1

Measurements and tests summary.


_jday_loc_serial_skcaodedsdtedhghprssluvzs
094Toronto#194calo3,menu,Must do HG now,Next schedule will be calo3 (now calo3 ZS= 93 ZS= 93.887),Waiting until ZA = -93 now -94.3,Waiting until ZA = -93 now -96.1,Waiting until ZA = -93 now -97.9,Waiting until ZA = -93 now -99.65153322121371312
(0 sec)

Relative and absolute humidity for day 2022094 (2 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
19414.12.2